OpenStudy (adelyn_helps):

The length of one leg of a right triangle is 21 ft and the length of the hypotenuse is 29 ft. What is the length of the other leg?

OpenStudy (adelyn_helps):

@ltrout

OpenStudy (steve816):

Use the Pythagorean theorem!

OpenStudy (mayankdevnani):



OpenStudy (mayankdevnani):

use pythagoras theorem ,

OpenStudy (ltrout):

This is Pythagorean Theorem. \[A^{2}+b^{2} =c^{2}\] . You have a leg and the hypotenuse. So 21^2 + B^2 = 29^2. Solve for B

OpenStudy (adelyn_helps):

I just learned this today so I have no idea what im doing! Please help me out here

OpenStudy (lyss):

I'll just restate that. 29^2 - 21^2 =

OpenStudy (ltrout):

First step would be to get b^2 on it's on side. So you could subtract either 21^2 or 29^2. Let's just go with B^2 = 29^2 - 21^2. Then you will need to find the values and then subtract them to get B^2. Then find the square root of that number and you're set.

OpenStudy (adelyn_helps):

so 20?

OpenStudy (mrnood):

@adelyn_helps first of all look at the triangle drawn above can oyu identify the right angle, and the hypotenuse? what is the length of the hypotenuse?

OpenStudy (ltrout):

Yes! 20 would be correct!

OpenStudy (lyss):

Yes. Sqrt of 400 is 20.

OpenStudy (mrnood):

A right angle triangle is made up of 3 sides - known as the hypotenuse and the 2 'legs' the hypotenuse is ALWAYS the longest side and is the side OPPOSITE the right angle

OpenStudy (mrnood):

no - it's \[29^2 = 21^2 + 20^2\]

OpenStudy (adelyn_helps):

Thannk you guys!

OpenStudy (ltrout):

Read it wrong, my bad @MrNood !

OpenStudy (adelyn_helps):

wait so 20 is incorrect or no?

OpenStudy (adelyn_helps):

Is 20 still correct @mrnood @ltrout

OpenStudy (ltrout):

Yes 20ft would the length of the final leg.

OpenStudy (adelyn_helps):

@ltrout I know you dont like being messaged but can you friend me I need to tell you something and then you can unfriend me again when im done!

OpenStudy (adelyn_helps):

thanks

OpenStudy (ltrout):

@adelyn_helps see if you're able to message now.